OpenStudy (anonymous):

There are two bags out of which Bag 1 contains 2 white and 3 red balls and bag 2 contains 4 white and 5 red balls. 1 ball is drawn at random from one of the bags and is found to be red. Find the probability that it was drawn from bag 2

OpenStudy (anonymous):

?how

OpenStudy (anonymous):

BAg 1 = 3/5 bag 2=5/9 i have calculate probabilty of red ball from bag 1 &2

OpenStudy (anonymous):

i didn't get u

OpenStudy (anonymous):

@Mertsj plse tell me

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

but there is no bag 3

OpenStudy (anonymous):

how many marbles there

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

excuse me

OpenStudy (anonymous):

This is a problem in conditional probability. The answer is not 1/2. I'll give 2 ways to look at this correctly. Then you can decide which is easier to foloow, but they both give the right answer, which again, is not 1/2.

OpenStudy (anonymous):

The first way is a less "scientific" and more of a common-sense approach. While all the marbles are still associated with a given bag, put a mark on each marble designating which bag it came from. Now you can dump them all out into one pile which correctly simulates the drawing from a bag at random. You are given that the marble is red, so separate out all the red marbles and look at them only. There are 8 in all. 5 have a mark on them that says "2". Your probability is 5/8. That's method one. Method 2 is more scientific and is in my next post.

OpenStudy (anonymous):

You are looking for P( 2 | R ) That is P ( 2 intersection R) / P(R) That is (5/14) / (8/14) and that equals 5/8 That is the second method and both give 5/8 To say that the probability is (incorrectly) 1/2 is like saying the probability of winning the lottery is 1/2, you either win or you lose as you are taking one of 2 possible outcomes.

OpenStudy (anonymous):

@tcarroll010 thnak u so much

OpenStudy (anonymous):

You are very welcome!

OpenStudy (anonymous):

The reason there is one helper showing an answer of 1/2 and another (me) showing the answer of 5/8 is that the problem ahs to be read carefully. What is being chosen at random? Is it the bag or is it the marble (ball)? If one uses a strict interpretation of the problem, the problem is saying that the ball is being chosen at random, not the bag. If the bag is what is being chosen at random, then the answer would indeed be 1/2. @msingh

OpenStudy (anonymous):

k